OpenStudy (anonymous):

Which transformations will produce similar, but not congruent, figures? Choose all answers that are correct.

OpenStudy (anonymous):

A. Square ABCD is rotated 270° clockwise and then dilated by a scale factor of to form square AꞌꞌBꞌꞌCꞌꞌDꞌꞌ. B. Square ABCD is reflected across the x-axis and then dilated by a scale factor of 2 to form square AꞌꞌBꞌꞌCꞌꞌDꞌꞌ. C. Square ABCD is dilated by a scale factor of and then translated 1 unit right to form square AꞌꞌBꞌꞌCꞌꞌDꞌꞌ. D. Square ABCD is translated 8 unit right and 8 units up and then reflected across the y-axis to form square AꞌꞌBꞌꞌCꞌꞌDꞌꞌ.

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

um a and idk

OpenStudy (mathmate):

The dilation factors in a and c are not legible. These factors are critical for this question.

OpenStudy (anonymous):

the one in c is \[\frac{ 1 }{ 3}\]

OpenStudy (mathmate):

When the dilation factor for a square is one, it will be both similar and congruent.

OpenStudy (mathmate):

when the dilation factor is different from one, transformed figure will be similar but not congruent.

OpenStudy (anonymous):

no the one in a is \[\frac{ 1 }{ 3}\] the one in c is \[\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

so then b and d you cant use fractions right?

OpenStudy (mathmate):

Rotations and translations preserve congruence unless there are other non-congruent transformations.

OpenStudy (mathmate):

Reflections also preserve congruence unless there are other non-congruent transformations.

OpenStudy (anonymous):

you lost me..

OpenStudy (anonymous):

so which ones could it be?

OpenStudy (mathmate):

Fractions are ok, as long as you can evaluate them. What's important for this question is whether the dilation factor equals one or different from one.

OpenStudy (mathmate):

Take the time to reread the question, and read my statements, and decide on which one(s) remain congruent, and which one(s) are similar but NOT congruent. There is more than one of the latter.

OpenStudy (anonymous):

whaaaa?

OpenStudy (anonymous):

mathmate was pretty confusing...

OpenStudy (mathmate):

...unless you read everything: "when the dilation factor is different from one, transformed figure will be similar but not congruent."